OpenStudy (anonymous):

Sheela buys 2 books for Rs. 182. She sells one at a loss of 5% and another at a profit of 8% . But she neither gains nor loses on the whole. Find the price of book which has sold at a profit -

OpenStudy (anonymous):

help me fast

OpenStudy (anonymous):

only tell me the answer

OpenStudy (anonymous):

zbay please help

OpenStudy (anonymous):

I think it's a trick question

OpenStudy (anonymous):

a) 112 b) 85 c) 70 d) 20

OpenStudy (anonymous):

112,70

OpenStudy (anonymous):

x+y=182 ---1 x*105/100+y92/100=182 x 105+y 92= 18200--2 eq1 x105+y105=182*105--3 eq3-eq1=13y=910 so y=910/13=70 and x=112

OpenStudy (anonymous):

sorry abdul the answer is only 70

OpenStudy (anonymous):

sorry abdul the answer is only 70

OpenStudy (anonymous):

but thanks

OpenStudy (anonymous):

soory me didnt read the qst properly